OpenStudy (jdm4iek):

which statement about the product is true? 5.643 x 4.66 irrational neither rational nor irrational cannot be determined rational

OpenStudy (jdm4iek):

irrational right? answer is 26.29638

OpenStudy (harman.singh):

Rational numbers are those which can be written as fractions. For eg. 1.5=3/2 Irrational numbers are those which cannot be written as fractions. For eg. 0.333 = ?? Does this help answer your question?

OpenStudy (518nad):

irrational has infinite decimals and it cannot be represented by an integer fraction

OpenStudy (jdm4iek):

hmm

OpenStudy (518nad):

this is rational

OpenStudy (518nad):

as there is a finite number of decimals , this can be represented as a fraction

OpenStudy (jdm4iek):

so what would 26.30 be in fraction form?

OpenStudy (harman.singh):

Do you think we can convert 26.29638 to a fraction?

OpenStudy (518nad):

for example 26.29638 =26296238/10000

OpenStudy (harman.singh):

Fraction form looks like this a/b . 26.30 is not in a fraction form

OpenStudy (518nad):

can u tell me what 0.30 is in fraction form

OpenStudy (jdm4iek):

wait

OpenStudy (jdm4iek):

so 1.5 is equal to 3/2

OpenStudy (518nad):

that is true

OpenStudy (jdm4iek):

but 0.333 is...

OpenStudy (vuriffy):

.333 is 1/3.

OpenStudy (518nad):

here okay let me help u understand decimals a bit

OpenStudy (jdm4iek):

yes

OpenStudy (518nad):

vurrify u r doing it again

OpenStudy (jdm4iek):

keep in mid it is almost 12 am im really tired

OpenStudy (vuriffy):

@518nad Just wanted to finish those dots. :)

OpenStudy (518nad):

><

OpenStudy (jdm4iek):

ok back on track peeps XD

OpenStudy (518nad):

okay let me help you understand decimals its much easier than u think okay

OpenStudy (518nad):

so take a look at any number lets say 123.12

OpenStudy (518nad):

every place of a number of telling u some unit of 10^n

OpenStudy (518nad):

for example in 123.45 you have 3 in 1s place 2 in 10s place 1 in 100s place 4 in 1/10s place 5 in 1/100s place

OpenStudy (518nad):

have you ever seen something like this before?

OpenStudy (jdm4iek):

yep

OpenStudy (518nad):

okay so any number you see can be written as a sum of these things

OpenStudy (518nad):

26.30 means you have 6 in 1s place 2 in 10s place 3in 1/10s place 0 in 1/100s place

OpenStudy (jdm4iek):

yeah

OpenStudy (518nad):

so u have 6*1 + 2*10 + 3 *1/10 + 0 * 1/100

OpenStudy (jdm4iek):

i understand decimals

OpenStudy (518nad):

can you add these fractions together

OpenStudy (jdm4iek):

but this is helping actually

OpenStudy (518nad):

xD

OpenStudy (518nad):

can you try to focus please

OpenStudy (518nad):

I am not misleading you, im telling you how to get to your solution

OpenStudy (jdm4iek):

so i got 26.3

OpenStudy (518nad):

26.30 2in 10s place 6 in 1s place 3 in 1/10s so you have 2 of 10, 6 of 1 and 3 of 1/10 2*10 + 6*1 +3 *1/10 = ??

OpenStudy (518nad):

do not use calculator use your fraction addition skills

OpenStudy (jdm4iek):

dangit

OpenStudy (jdm4iek):

2x10=20 6x1= 6 3x1/10 = 0.3 26..3

OpenStudy (518nad):

keep it as a fraction buddy throw away decimals

OpenStudy (518nad):

3*1/10 = 3/10

OpenStudy (518nad):

what is 26 + 3/10

OpenStudy (518nad):

26/1 + 3/10

OpenStudy (518nad):

we need common denominator to add fractions

OpenStudy (jdm4iek):

ok

OpenStudy (518nad):

can you find this 26/1 = ___/10 what goes in the blank

OpenStudy (jdm4iek):

3?

OpenStudy (518nad):

nono i just wanna rewrite 26/1 as a denominator with 10 now

OpenStudy (518nad):

so that we can add it to 3/10

OpenStudy (jdm4iek):

so should we divide 26 / 10?

OpenStudy (518nad):

that is no longer the same as 26/1 though is it

OpenStudy (518nad):

look we want to add these 2 26/1 + 3/10 but they dont have same denominator we need to express 26 in another way with 10 as denominator

OpenStudy (518nad):

260/10 = 26/1

OpenStudy (jdm4iek):

oooohh

OpenStudy (jdm4iek):

i feel dumb :/ keep goin though

OpenStudy (518nad):

mhm now 26/1 + 3/10 =260/10+3/10 =263/10

OpenStudy (518nad):

its okay you just gotta understand fractions

OpenStudy (jdm4iek):

now we have 263/10

OpenStudy (518nad):

right which is 26.3

OpenStudy (jdm4iek):

yep

OpenStudy (518nad):

-.-

OpenStudy (jdm4iek):

what?

OpenStudy (518nad):

i am so not happy with this rn... u really need to learn fractions

OpenStudy (518nad):

i dont even know if u realize you have the answer there

OpenStudy (jdm4iek):

i have really no idea how to do fractions my teachers never really tought me

OpenStudy (518nad):

okay ill help you out a bit fractions are really nice to learn, it will make math fun and easy

OpenStudy (jdm4iek):

ok youre the boss

OpenStudy (518nad):

okay you need to understand 2 or 3 main things to really get fractions

OpenStudy (jdm4iek):

alright

OpenStudy (518nad):

what a whole is, and what it makes to take parts of wholes, and how we can think about adding parts. 2nd is understanding fraction to decimal conversion

OpenStudy (518nad):

fractions when the denominator is some power of 10 can rewritten as a decimal

OpenStudy (518nad):

and you need to understand what equivalent fractions are and why they are equal and thats pretty much everything

OpenStudy (jdm4iek):

ok

OpenStudy (518nad):